Question 4 5 pts Consider the equation nº X () + X(T) - 0 with the following boundary conditions. X(0) + X(0) -0, and X(n)

n=5
Answer #1

Case I so, Given that ñ x²(x) + Xx (2) = X(k) + W x (x) =0, when a=0, Then, X (1) = the general salh X (W) ciute and X(W)case 3 then when ao X (W) + 7 X (K) =0. the goneral Sain X (2) Gus an tosin on X(x) = 1 sain 2x + 9 7 6 24 now using the bo